Martinez Inc. issued $4,320,000 of 9%, 10-year convertible bonds on June 1, 2020, at 99 plus accrued interest. The bonds were dated April 1, 2020, with interest payable April 1 and October 1. Bond discount is amortized semiannually on a straight-line basis.

On April 1, 2021, $1,620,000 of these bonds were converted into 25,000 shares of $20 par value common stock. Accrued interest was paid in cash at the time of conversion.

(a) Prepare the entry to record the interest expense at October 1, 2020. Assume that accrued interest payable was credited when the bonds were issued.
(b) Prepare the entry to record the conversion on April 1, 2021. (Book value method is used.) Assume that the entry to record amortization of the bond discount and interest payment has been made.


(Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Round answers to 0 decimal places, e.g. 5,125.)

No.

Account Titles and Explanation

Debit

Credit

(a)

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

(b)

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount
Answer #1

Total Discount = $ 4320 000 - ($1320000*0.99) = $ 00 $4276800 $ 4320000 43200 ( remaining month cox22-2) Discant amastized 432 (6) un mortized band discount = (43200437/[u3200 x 3 %*) 3x10 118 = 16200 - 1373 = $14827 Journal entry : Account title and